solve the equation algebraically.
4x - 3 = 2
&
9 = 10 - x

Answer by ewatrrr(24785)


Hi,
4x - 3 = 2
4x = 5
x = (5/4)
9 = 10 - x
9 + x = 10 -x + x
9 - 9 + x = 10 - 9
x = 1